# From Coarse to Fine-Grained Emotion Annotation: An Immediate Recall Paradigm with Validation through Physiological Evidence and Recognition Performance

### v1

@@ VI Conclusion
To address the prevalent challenges of label noise in video-induced emotion datasets, this paper introduces FIRMED, a multimodal dataset developed using an immediate recall-based annotation paradigm. Our comprehensive validation demonstrates the paradigm’s methodological rigor and empirical effectiveness across multiple analytical dimensions.
Physiologically, we established temporal alignment between participant-reported emotional events and objective neurophysiological signatures, including emotion-specific oscillatory patterns across EEG frequency bands and distinctive autonomic activation profiles. This provides validation of the subjective-objective correspondence that has long been theoretically postulated but rarely systematically verified in emotion research. Computationally, classification models trained on precisely annotated data demonstrated consistent and substantial performance improvements across eight distinct algorithms, and multimodal approaches demonstrated superior performance compared to EEG-only configurations. These computational findings align with conclusions from SEED-VII [21], reinforcing that annotation quality serves as a critical determinant of recognition performance.
We acknowledge several limitations that warrant consideration. The dataset scale constrains statistical power and generalizability across diverse populations. The fixed 4-second analysis window, while physiologically informed, may not optimally capture the heterogeneous temporal dynamics of emotional responses across individuals and emotional contexts. Future investigations should prioritize expanding participant diversity across demographics, implementing adaptive temporal windowing techniques that accommodate individual response variability.

### v2

@@ VI Conclusion
To address the label noise and temporal imprecision common in traditional whole-trial annotation methods in video-induce emotion datasets, this study proposed and validated a novel immediate recall paradigm, inspired by the Peak-End Rule. This paradigm leverages an immediate video replay phase, enabling participants to accurately identify and label discrete timestamps corresponding to the subjectively experienced peak of emotion. Guided by fresh recollection, the paradigm minimizes reliance on biased memory reconstruction. The FIRMED dataset was constructed based on this paradigm.
Comprehensive validation confirmed the paradigm’s effectiveness. Physiological validation revealed strong temporal correspondence between subjective marked timestamps and objective, emotion-specific EEG neural patterns and GSR-SCR phenomena of autonomic arousal, confirming the capture of meaningful physiological events. Comparative experiments showed reduced annotation uncertainty compared to delayed recall. Critically, computational validation demonstrated that classifiers trained on FIRMED significantly outperformed those trained on traditional whole-trial labels across multiple algorithms and modalities. This work robustly demonstrates that for physiological emotion recognition, annotation precision, rather than mere data scale, is a critical factor in determining model performance.
We acknowledge that future work could improve upon our methodology. For instance, designing a counterbalanced or randomized stimulus sequence would be a methodological improvement over our fixed-order design. Furthermore, exploring adaptive windowing strategies to model varying emotional latencies remains a valuable direction for future research.